A man borrows ₱6,400 from a loan association. In repaying this debt he has to pay ₱400 at the end of every 3 months on the principal and a simple interest of 16% on the principal outstanding at that time. Determine the total amount he has paid after paying all his debt.


1
Expert's answer
2021-10-14T10:49:38-0400

Solution:

The correct answer is ₱8,576.00.


"\frac{3}{12} = 0.25"


Interest 1 = 6400 "\times" 0.25 "\times" 0.16 = 256

Interest 2 = 6000 "\times" 0.25 "\times" 0.16 = 240

Interest 16 = 400 "\times" 0.25 "\times" 0.16 = 16

Total Principal = 6400 + S

S = "\frac{n}{2} \times" (a1 +an) = "\frac{16}{2} \times (256 + 16) = 2,176"


TP = 6400 + 2176 = 8576